OpenStudy (anonymous):

if f(x)=(1/2)x-6. what is the equation for f^-1(x)? ...2x-6 ...2x+6 ...2(x-6) ...2(x-6)

OpenStudy (anonymous):

this one is doable. put \[y=\frac{1}{2}x-6\]switch x and y \[x=\frac{1}{2}y-6\] solve for y \[x+6=\frac{1}{2}y\] \[2(x+6)=y\]

OpenStudy (anonymous):

must be a typo on your answers. you have \[2(x-6)\] twice. perhaps one is \[2(x+6)\]?

OpenStudy (anonymous):

yes one is +6

OpenStudy (anonymous):

that is the one!

OpenStudy (anonymous):

f(x)=2x+12

OpenStudy (anonymous):

why do you think it's +12??? when it can be simplified more

OpenStudy (anonymous):

yeah i'm just distributing it. you can have also y=2(x+6)

OpenStudy (anonymous):

yeah i got it

OpenStudy (anonymous):

thank you